# ITDR: An Instruction Tuning Dataset for Enhancing Large Language Models in Recommendations

### v1

@@ 4.3. Results and Analysis of Key Research Questions
$\bullet$ After fine-tuning with ITDR, the performance disparities among the four models were significantly reduced. Taking the UII task as an example, the pUII metric gap between the un-tuned LLaMA-3.2-3B and Qwen2.5-7B exceeded 10, while after fine-tuning, this gap was reduced to within 2. These empirical results demonstrate that ITDR exhibits broad model adaptability and can effectively enhance the performance of different types of LLMs on recommendation tasks.
RQ2: Do potential interactions exist between different root tasks?
@@
(a) Average performance of UII.
(b) Average performance of UIU.
Figure 3. Average performance of removing different root tasks.
@@
In this section, we investigate the synergistic mechanisms among subtasks through task-level ablation experiments. Specifically, we sequentially remove specific subtasks (e.g., the Top-k recommendation task) from ITDR, then fine-tune GLM-4-9B using the remaining training data, and evaluate its performance on the complete test set.
@@
(a) Ablation studies of root tasks in removing task descriptions during fine-tuning.
(b) Average performance of using different data volumes for fine-tuning.
Figure 4. Performance of using task description and different data volumes.
ITDR provides detailed task descriptions for each subtask, a design intended to enhance the model’s task comprehension and enrich data diversity. In this section, we focus on evaluating the practical impact of these task descriptions on model performance. By removing the task description component from the instructions and fine-tuning the GLM-4-9B model, we obtained the experimental results illustrated in Figure 4(a) and Table 2. Based on these findings, we draw the following conclusions:
@@
In this section, we investigate the impact of fine-tuning data scale on model performance. Using stratified sampling, we created three subsets of training data at 25%, 50%, and 75% proportions to fine-tune the GLM-4-9B. The model’s performance was then evaluated on the test set, with experimental results shown in Figure 4(b). Our findings reveal that:
@@
$\bullet$ The fine-tuned models demonstrate superior performance on both UII and UIU tasks compared to the baseline model with larger parameter sizes. Notably, benefiting from the advantage of massive parameter scale, larger models still exhibits significantly better performance on the UII task than untuned open-source models. Further analysis reveals distinct task preference characteristics among different models: GPT-3.5-Turbo performs more effectively on UIU tasks, while Gemini 2.5 Flash demonstrates stronger processing capabilities on UII tasks, indicating fundamental differences in how different model architectures understand and solve tasks. This series of findings reaffirms the significant value of our developed ITDR dataset in recommendation tasks.
@@ 5. Conclusion
In this paper, we propose ITDR, an instruction tuning dataset constructed from two dimensions—user-item interaction and user-item understanding—effectively filling the gap of natural language datasets in recommendation systems. We systematically investigate leveraging ITDR to enhance LLMs’ performance on recommendation tasks. Through extensive experiments, we explore several key aspects, including task relationships, task description design, and the impact of data volume in instruction tuning, demonstrating ITDR’s effectiveness in enhancing LLM-based recommendations. We expect ITDR to advance research on LLMs for recommendation and stimulate further academic exploration of instruction tuning methodologies.
@@
In this work, we construct a sufficient dataset for instruction tuning on recommendation systems and trained fine-tuned models based on this dataset. We identify the limitations of the current research and proposed directions for future improvements.
In terms of the dataset, although we integrat 7 types of recommendation subtasks from 13 source datasets, a large number of datasets and tasks remain unincorporated. Experimental results indicate that introducing more diverse data could further enhance the coverage and diversity of the dataset, which may positively impact model performance.
Regarding model parameters, our study only validat the effectiveness of models with parameters ranging from 3B to 9B. Future work could explore the impact of instruction tuning on larger-scale models (e.g., 13B, 70B-level) and conduct comparisons with existing closed-source LLMs, which would hold significant research value.
For the evaluation methods, due to resource constraints, we primarily focuse on average performance metrics at the task level. Subsequent research could analyze the interaction effects between different datasets at a finer granularity, providing more instructive recommendations for optimizing the composition of instruction tuning datasets.
@@ Acknowledgements.
This work was supported in part by the National Key Research and Development Program of China under Grant (2023YFC3310700) and the National Natural Science Foundation of China (62202041).
Appendix A Details about Tasks and Datasets
A.1. User-Item Interaction
In recommendation systems, user-item interaction data serves as the core element for system optimization, capturing the multidimensional behavioral trajectories between users and recommended content. This includes explicit feedback (such as ratings and purchases) and implicit signals (such as clicks and browsing duration). These interaction data inherently reflect the complex mapping relationships between user preferences and item features, providing the foundational data for building accurate personalized recommendation models. Among these, how to effectively model user-item interaction relationships not only directly impacts the accuracy of recommendation results but also significantly influences user satisfaction. To this end, we design a series of subtasks focused on user-item interactions, aiming to enhance the capability of LLMs in capturing user behaviors and item features, with a particular emphasis on evaluating interaction modeling tasks across the following four tasks.
Rating Prediction (RP)
Top-K Recommendation (TKR)
Cross-Domain Recommendation (CDR)
Table 3. Selection strategies for source and target domains in cross-domain recommendation task.
Source Target
Beauty Fashion
CDs & Vinyl Video Games
Appliances Industrial & Scientific
Home & Kitchen Grocery & Gourmet Food
Next Item Recommendation (NIR)
A.2. User-Item Understanding
In recommendation systems, user-item understanding serves as the core dimension of system cognition. By deeply mining the bidirectional semantic associations between user characteristics and item attributes, it provides critical support for personalized recommendations. This process involves in-depth analysis at two levels: On the user side, the system constructs precise user profiles by analyzing historical behavioral data; on the item side, it focuses on analyzing target audience characteristics to form comprehensive item profiles. These multidimensional understandings establish mapping relationships between users and items, not only improving recommendation accuracy but also laying the cognitive foundation for explainable recommendation logic. Based on this, we design three user-item interaction subtasks aimed at enhancing the large language model’s joint reasoning and matching capabilities for user attributes and item features.
User Attribute Prediction (UAP)
Interest Recognition (IR)
The primary objective of interest recognition is to analyze user‘s historical interaction data to identify and summarize their preferences, thereby enabling more accurate personalized services. For this subtask, we select the following datasets as data sources: Amazon Books Reviews2, Anime Dataset 20231, and Steam (steam).
Target User Identification (TUI)
Appendix B Statistics of ITDR and Additional Results
@@
According to the experimental data in Table 6 – Table 8, the fine-tuned model shows significant improvements in performance metrics across all tasks, particularly in sequence recommendation tasks (TKR, CDR, NIR) and generation tasks (IR, TUI). The results fully validate the effectiveness of ITDR in improving the performance of LLMs in recommendation tasks.
@@
Appendix C Data Examples
Data examples for each subtask are presented in Table 18 – Table 24. The construction process for each instruction is shown below.
@@ Appendix C Data Examples
When constructing instructions for the CDR task based on the Amazon Review 2023 dataset, we implement the following process: first, we extract the user’s first 10 interaction records sorted in ascending chronological order from the source domain as prior knowledge input for the model; second, we select the top 5 positive sample records from the target domain, sorted in ascending order by time, and combine them with 5 randomly sampled negative records to form the candidate ranking set. The model’s expected output is the correct sorted order of these samples. Table 20 shows the details about the instruction.
@@
Taking the MovieLens 1M dataset as an example, we designed an instruction data construction method for the UAP task: First, we filtered out each user’s movie-watching records with ratings above 3 as positive samples and sorted them in ascending order by viewing time. Then, we used the user’s static attribute features as the target variables for the model to predict. As shown in Table 22, we provide specific instruction examples to illustrate the data construction process.
@@
Table 24 demonstrates an example of constructing TUI task instructions based on the Steam dataset. This construction process includes the following steps: First, for each game, extract its key attributes, including feature information such as the game’s name, genre, tags, and gameplay. Second, randomly sample real player reviews of the game (which may contain negative feedback). Similar to IR tasks, we use DeepSeek-V3 to generate standard answer data as the expected output for the instructions.
Appendix D Notations and Terminologies
We summarize the key notations used in this paper in Table 4.
Table 4. Key notations.
Symbol Description
UII User-Item Interaction
UIU User-Item Understanding
RP Rating Prediction
TKR Top-K Recommendation
CDR Cross-Domain Recommendation
NIR Next Item Recommendation
UAP User Attribute Prediction
IR Interest Recognition
TUI Target User Identification
$ST$ set of subtasks
$M$ set of evaluation metrics for sequential recommendation tasks
$D_{st}$ set of all subtasks data
$pUII$ performance of user-item interaction
$pUIU$ performance of user-item understanding
$D_{UAP}$ set of UAP task
$M^{\prime}$ set of evaluation metrics for IR and TUI tasks
@@ Appendix D Notations and Terminologies
Top-K Recommendation Amazon Reviews 2023
N@10
H@1
P@5
5,276 435.32 264.53
@@
Cross-Domain Recommendation Amazon Reviews 2023
N@10
H@1
P@5
4,118 651.50 313.67
@@
User Attribute Prediction Last.FM 1K
ACC(country)
ACC(gender)
RMSE(age)
943 590.56 11.44
User Attribute Prediction Last.FM 360K
ACC(country)
ACC(gender)
RMSE(age)
6,250 267.11 12.72
User Attribute Prediction Anime Dataset 2023
ACC(location)
ACC(gender)
RMSE(Birth year)
6,250 571.91 16.55
User Attribute Prediction BookCrossing
ACC(location)
RMSE(age)
6,250 71.81 14.30
User Attribute Prediction MovieLens 1M
ACC(age group)
ACC(gender)
ACC(occupation)
6,040 310.24 18.30
Interest Recognition Amazon Books Reviews
BLEU-1&2,
ROUGE-1&2&L
6,250 1,675.46 107.18
@@
Target User Identification Amazon Reviews 2023
BLEU-1&2
ROUGE-1&2&L
6,300 370.89 253.21
@@
Table 7. Performance on the user-item understanding task.

### v2

@@ 4.3. Results and Analysis of Key Research Questions
$\bullet$ After fine-tuning with ITDR, the performance disparities among the four models were significantly reduced. Taking the UII task as an example, the pUII metric gap between the un-tuned LLaMA-3.2-3B and Qwen2.5-7B exceeded 10, while after fine-tuning, this gap was reduced to within 2. These results demonstrate that ITDR exhibits broad model adaptability and can effectively enhance the performance of different types of LLMs on recommendation tasks.
RQ2: Do potential interactions exist between different root tasks?
@@
(a) Average performance of UII.
(b) Average performance of UIU.
Figure 3. Average performance of removing different root tasks.
@@
Figure 4. Ablation studies of using no task descriptions during fine-tuning.
@@
In this section, we investigate the synergistic mechanisms among subtasks through task-level ablation experiments. Specifically, we sequentially remove subtasks (e.g., the Top-k recommendation task) from ITDR, then fine-tune GLM-4-9B using the remaining training data, and evaluate its performance on the complete test set.
@@
ITDR provides detailed task descriptions for each subtask, a design intended to enhance the model’s task comprehension and enrich data diversity. In this section, we focus on evaluating the practical impact of these task descriptions on model performance. By removing the task description component from the instructions and fine-tuning the GLM-4-9B model, we obtained the experimental results illustrated in Figure 4 and Table 4. Based on these findings, we draw the following conclusions:
Figure 5. Average performance of using different data volumes for fine-tuning.
@@
In this section, we investigate the impact of fine-tuning data scale on model performance. Using stratified sampling, we created three subsets of training data at 25%, 50%, and 75% proportions to fine-tune the GLM-4-9B. The model’s performance was then evaluated on the test set, with experimental results shown in Figure 5. Our findings reveal that:
@@
$\bullet$ The fine-tuned models demonstrate superior performance on both UII and UIU tasks compared to the baseline model with larger parameter sizes. Notably, benefiting from the advantage of massive parameter scale, larger models still exhibit significantly better performance on the UII task than untuned open-source models. Further analysis reveals distinct task preference characteristics among different models: GPT-3.5-Turbo performs more effectively on UIU tasks, while Gemini 2.5 Flash demonstrates stronger processing capabilities on UII tasks, indicating fundamental differences in how different model architectures understand and solve tasks. This series of findings reaffirms the significant value of our developed ITDR dataset in recommendation tasks.
@@ 5. Conclusion
In this paper, we propose ITDR, an instruction tuning dataset constructed from two dimensions–user-item interaction and user-item understanding–effectively filling the gap of natural language datasets in recommendation systems. We systematically investigate leveraging ITDR to enhance LLMs’ performance on recommendation tasks. Through extensive experiments, we explore several key aspects, including task relationships, task description design, and the impact of data volume in instruction tuning, demonstrating ITDR’s effectiveness in enhancing LLM-based recommendations. We expect ITDR to advance research on LLMs for recommendation and stimulate further academic exploration of instruction tuning methodologies.
@@
In this work, we construct a dataset for instruction tuning on recommendation systems and trained fine-tuned models based on this dataset. We identify the limitations of the current research and proposed directions for future improvements.
In terms of the dataset, although we have integrated seven distinct recommendation subtasks across 13 source datasets, this represents only a fraction of the available data landscape, leaving numerous potential datasets and specialized tasks untapped. Our experimental findings suggest that systematically expanding this scope to include more diverse data would not only broaden the dataset’s coverage but also significantly enhance its semantic richness, thereby unlocking further gains in model performance and generalization.
Regarding model parameters, our study only validate the effectiveness of models with parameters ranging from 3B to 9B. Future work could explore the impact of instruction tuning on larger-scale models (e.g., 13B, 70B-level) and conduct comparisons with existing closed-source LLMs, which would hold significant research value.
For the evaluation methods, due to resource constraints, we primarily focused on evaluating the overall average performance across tasks. Future work could further dissect the links and transfer effects among different datasets at a finer level, thereby providing a clearer view of how data distribution characteristics affect the emerging recommendation abilities of LLMs. Such findings would offer solid guidance for optimizing data composition and sampling strategies during instruction tuning of LLMs for recommendation tasks.
@@ Acknowledgements.
This work was supported in part by the National Key Research and Development Program of China under Grant (2023YFC3310700), the National Natural Science Foundation of China (62572040, 62202041), the Beijing Natural Science Foundation (JQ24019), and the Fundamental Research Funds for the Central Universities (2025JBMC011).
Appendix A Statistics of ITDR and Additional Results
@@
According to the experimental data in Table 6 – Table 8, the fine-tuned models show significant improvements in performance metrics across all tasks, particularly in sequence recommendation tasks (TKR, CDR, NIR) and generation tasks (IR, TUI). Taking the performance of GLM-4 before and after fine-tuning as an example, on the MicroLens dataset under the TKR task, the H@1 metric increases from 0.5520 to 0.6960; on the PixelRec dataset for the TUI task, the ROUGE-1 score also rises from 0.3359 to 0.5246. These findings strongly validate the effectiveness of ITDR in enhancing the recommendation capabilities of LLMs.
@@
Appendix B Data Examples
Data examples for each subtask are presented in Table 18 – Table 24.
@@ Appendix B Data Examples
When constructing instructions for the CDR task based on the Amazon Review 2023 dataset, we first use the user’s earliest 10 interaction records from the source domain (sorted chronologically in ascending order) as prior knowledge input. Second, we select the earliest 5 positive samples from the target domain (also sorted chronologically) and combine them with 5 randomly sampled negative samples to form a candidate ranking set. The model is expected to output the correct ranking order of these samples. See Table 20 for details.
@@
Taking the MovieLens 1M dataset as an example, we designed an instruction data construction method for the UAP task: First, we filtered out each user’s movie-watching records with ratings above 3 as positive samples and sorted them in ascending order by viewing time. Then, we used the user’s static attribute features as the target variables for the model to predict. As shown in Table 22, we provide an example to illustrate the data construction process.
@@
Table 24 demonstrates an example for constructing TUI task instructions based on the Steam dataset. First, for each game, extract its key attributes, including feature information such as the game’s name, genre, tags, and gameplay. Second, randomly sample real player reviews of the game (which may contain negative feedback). Similar to IR tasks, we use DeepSeek-V3 to generate standard answer data as the expected output for the instructions.
@@
Top-K Recommendation Amazon Reviews 2023 N@10 H@1 P@5 5,276 435.32 264.53
@@
Cross-Domain Recommendation Amazon Reviews 2023 N@10 H@1 P@5 4,118 651.50 313.67
@@
User Attribute Prediction Last.FM 1K ACC(country) ACC(gender) RMSE(age) 943 590.56 11.44
User Attribute Prediction Last.FM 360K ACC(country) ACC(gender) RMSE(age) 6,250 267.11 12.72
User Attribute Prediction Anime Dataset 2023 ACC(location) ACC(gender) RMSE(Birth year) 6,250 571.91 16.55
User Attribute Prediction BookCrossing ACC(location) RMSE(age) 6,250 71.81 14.30
User Attribute Prediction MovieLens 1M ACC(age group) ACC(gender) ACC(occupation) 6,040 310.24 18.30
Interest Recognition Amazon Books Reviews BLEU-1&2, ROUGE-1&2&L 6,250 1,675.46 107.18
@@
Target User Identification Amazon Reviews 2023 BLEU-1&2 ROUGE-1&2&L 6,300 370.89 253.21
@@
Table 7. Performance on the user-item understanding tasks.